OpenStudy (anonymous):

If x*x + 2x - 35 = 0, then x = ?

OpenStudy (shamil98):

x^2 + 2x - 35 =0 a = 1 b= 2 c = -35 Use the quadratic formula for this, \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (wolf1728):

Or by factoring (X - 5)(x + 7)

OpenStudy (shamil98):

Yes, that works as well, if you know how to factor, that is.